OpenStudy (anonymous):

(-2, -3) and (-6, -5) the equation of the line is slope-intercept form is y=

OpenStudy (anonymous):

So you first need to find the slope (m) of your equation. M= (y2-y1)/(x2-x1) Now it does not matter which y coordinate you assign as y2 or y1 just as long as the x-coordinate coincides with the y coordinate you choose. So let y2=-5 (then x2=-6), and let y1=-3 (so x1=-2) Then the slope equation will look like this: m=(-5-(-3))/(-6-(-2))=(-2)/(-4)=1/2 So m = 1/2. Now you can use what bahrom7893 showed you: y-y1 = M(x-x1) y - (-3) = (1/2)(x-(-2)) y + 3 = (1/2)x - (1) y = (1/2)x-1-3 y = (1/2)x-4 Does that make sense?